OpenStudy (abbey2804):

can someone help me.... use the quadratic formula to solve 2x squared-3x-1=0

pooja195 (pooja195):

Hi there! Welcome to OpenStudy :-) First identify the abc values then plug them in :) \(\normalsize\color{ slate }{{\bbox[5pt, lightcyan ,border:2px solid black ]{ \LARGE{x=~} \huge{ \frac{-\color{magenta}{b} \pm\sqrt{ \color{magenta}{~b} ^2-4 \color{blue}{a} \color{red}{c}~}}{2 \color{blue}{a}} }~ }}}\) when the equation is \(\large\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\).

OpenStudy (abbey2804):

okay, thank you!!

pooja195 (pooja195):

You're welcome! :D

OpenStudy (robtobey2):

x = 1/4 (3 \[PlusMinus] Sqrt[17]) Cell[BoxData[ RowBox[{"x", "\[Rule]", RowBox[{ FractionBox["1", "4"], " ", RowBox[{"(", RowBox[{"3", "\[PlusMinus]", SqrtBox["17"]}], ")"}]}]}]], "Input", CellChangeTimes->{{3.677903902754315*^9, 3.6779039028315573`*^9}}]